OpenStudy (mathhelp346):

Deternine whether each proportion is true. Show how. 1. 15/9 = 35/21 2. 56/24 = 49/21

OpenStudy (lukecrayonz):

divide them.

OpenStudy (anonymous):

Simplify them down to lowest terms

OpenStudy (anonymous):

\[1. \space \frac{5}{3} = \frac{5}{3}\] so number one is true

OpenStudy (anonymous):

1. true 2. 7/3 = 7/3 so, they are true

OpenStudy (anonymous):

\[2. \space \frac{7}{3} = \frac{7}{3}\] 2 is also true